# UDC-VIT: A Real-World Video Dataset for Under-Display Cameras

### v1

@@ Abstract
Under Display Camera (UDC) is an advanced imaging system that places a digital camera lens underneath a display panel, effectively concealing the camera. However, the display panel significantly degrades captured images or videos, introducing low transmittance, blur, noise, and flare issues. Tackling such issues is challenging because of the complex degradation of UDCs, including diverse flare patterns. Despite extensive research on UDC images and their restoration models, studies on videos have yet to be significantly explored. While two UDC video datasets exist, they primarily focus on unrealistic or synthetic UDC degradation rather than real-world UDC degradation. In this paper, we propose a real-world UDC video dataset called UDC-VIT. Unlike existing datasets, only UDC-VIT exclusively includes human motions that target facial recognition. We propose a video-capturing system to simultaneously acquire non-degraded and UDC-degraded videos of the same scene. Then, we align a pair of captured videos frame by frame, using discrete Fourier transform (DFT). We compare UDC-VIT with six representative UDC still image datasets and two existing UDC video datasets. Using six deep-learning models, we compare UDC-VIT and an existing synthetic UDC video dataset. The results indicate the ineffectiveness of models trained on earlier synthetic UDC video datasets, as they do not reflect the actual characteristics of UDC-degraded videos. We also demonstrate the importance of effective UDC restoration by evaluating face recognition accuracy concerning PSNR, SSIM, and LPIPS scores. UDC-VIT enables further exploration in the UDC video restoration and offers better insights into the challenge. UDC-VIT is available at our project site.
@@ 1 Introduction
An under-display camera (UDC) is an imaging system where the camera is positioned beneath the display. Modern smartphones, including the Samsung Galaxy Z-Fold series (Samsung Electronics Co., Ltd., 2021; 2022; 2023; 2024) and the ZTE Axon series (ZTE Corporation, 2020; 2021; 2022) have adopted UDCs. The UDC area, depicted in Figure 1, serves as display space under normal circumstances and acts as the light’s passage to the camera when capturing pictures or videos. This design allows for a larger screen-to-body ratio, meeting the common consumer demand for a full-screen display without a camera hole or notch. However, UDC introduces severe and complex image degradations such as reduced transmittance, noise, blur, and flare in a single image or video frame. Moreover, motion is also involved in UDC videos.
@@
The degradation in UDC arises from the diffraction of incoming light by the display pixels at a micrometer scale (Qin et al., 2016). Modern UDC smartphones have lower pixel density in the UDC area to minimize this diffraction, as described in Figure 1(c). Since a lower pixel density prevents natural video viewing, improving the video quality captured by the UDC is essential.
@@
Many studies have investigated UDC image datasets. These include synthetic datasets like T-OLED/P-OLED (Zhou et al., 2021) and SYNTH (Feng et al., 2021). Additionally, there exists a pseudo-real dataset (Feng et al., 2023) and a real-world dataset called UDC-SIT (Ahn et al., 2024).
@@
Ahn et al. (2024) demonstrate the importance of training DNN models using a real-world UDC dataset because the synthetic UDC datasets do not reflect the actual characteristics of UDC-degraded images. However, a real-world UDC video dataset and restoration model have yet to be introduced. Although several studies address the synthetic UDC video datasets (Chen et al., 2023; Liu et al., 2024), they have some limitations because they do not completely reflect the properties of actual UDC videos. There are two main challenges in constructing a real-world UDC video dataset. One is to find a matching pair of the UDC-distorted and ground-truth videos with high alignment accuracy. The other is to synchronize the time for all frames when capturing videos.
@@
This paper proposes a new UDC video dataset called UDC-VIT (UDC’s VIdeo by Thunder Research Group). As far as we know, it is the first real-world UDC video dataset to overcome the problems of the existing UDC video datasets.
Using a non-polarizing cube beam splitter (Thorlabs, 2015), we create a video-capturing system to minimize discrepancies between paired frames. We cut the UDC area of a smartphone display (e.g., Samsung Galaxy Z-Fold 5 (Samsung Electronics Co., Ltd., 2023)) and attach it to the beam splitter. Two Arducam Hawk-Eye (IMX686) camera modules (Arducam, 2022) are placed on both sides of the beam splitter. These modules, operated by a Raspberry Pi 5 (Arducam, 2023), capture synchronized video frame pairs using the Message Passing Interface (MPI) barrier.
Figure 2 shows our UDC video capturing system. Despite the meticulous design, inevitable pixel-position difference occurs. We correct this difference between the two matched frames for the same scene by using the DFT (Brigham, 1988) following the previous work by Ahn et al. (2024).
The contributions of this paper are summarized as follows:
  • We address the limitations of existing datasets, including unrealistic degradations, flares, and artifacts, emphasizing the need for a real-world dataset. (Section 4 and Appendix B)
  • We provide UDC-VIT, a real-world UDC video dataset that accurately reflects actual UDC degradations, ensuring precise spatial and temporal alignment through our meticulously designed video-capturing system (Section 3 and Appendix A).
  • We describe UDC-VIT’s effectiveness through extensive experiments, comparing it with an existing synthetic dataset using six deep-learning models. High-quality datasets are crucial for advancing representation learning. (Section 5, Appendix C and D)
  • We highlight the importance of restoring UDC degradation for applications like Face ID by evaluating face recognition accuracy at different restoration levels. Our dataset includes real-world face images, enhancing its relevance for practical use. (Section 5)
@@ 2 Related Work
There has been extensive research on UDC still image datasets. Zhou et al. (2021) propose the T-OLED/P-OLED datasets. Images are displayed on a monitor, and paired images are captured with and without a T-OLED/P-OLED display in front of the camera. However, due to the limited dynamic range of the monitor, flares are almost absent in their datasets. Feng et al. (2021) propose the SYNTH dataset. They convolve the measured point spread function (PSF) of ZTE Axon 20 (ZTE Corporation, 2020) with clean images (Haven, 2020), exhibiting flares. However, it has limitations such as the absence of noise and spatially variant flares. Notably, UDC distortion gradually increases from the center of the camera lens to outwards, leading to spatially distorted flares (Yoo et al., 2022). Feng et al. (2023) propose a pseudo-real dataset by capturing paired images of similar scenes using two cameras (e.g., ZTE Axon 20 UDC (ZTE Corporation, 2020) and iPhone 13 Pro camera (Apple Inc., 2021)). However, they use two cameras, leading to geometric misalignment. They improve the geometric misalignment using AlignFormer (Feng et al., 2023). Nonetheless, they encounter challenges with alignment accuracy. Ahn et al. (2024) propose a real-world dataset called UDC-SIT and an image-capturing system. They attach Samsung Galaxy Z-Fold 3 (Samsung Electronics Co., Ltd., 2021)’s UDC area to a lid. Paired images are acquired by opening and closing the lid onto the Samsung Galaxy Note 10’s standard camera (Samsung Electronics Co., Ltd., 2019). They use DFT to align the misalignment between the paired images that occurs during the opening and closing of the lid. The images in the UDC-SIT dataset contain the actual UDC degradation (e.g., spatially variant flares). Finally, Wang et al. (2024) and Tan et al. (2023) propose face datasets for facial expression recognition or restoration. They are synthesized using a GAN-based model trained on the T/P-OLED dataset (Zhou et al., 2021), which lacks realistic UDC degradation, particularly flares. Moreover, the datasets are not publicly available.
@@
Research has been conducted on synthetic UDC video datasets. Chen et al. (2023) propose the PexelsUDC-T/P dataset. They train a GAN-based UDC video generation model using T-OLED/P-OLED datasets (Zhou et al., 2021), which do not show flares. They generate UDC-degraded videos using clean videos (Pexels, 2014). Moreover, the datasets are not publicly available. Liu et al. (2024) propose the VidUDC33K dataset. They convolve the measured PSF on the clean video frames (Haven, 2020) to show flares. They simulate the dynamic change of the PSF (Kwon et al., 2021) between consecutive frames following the previous work (Babbar & Bajaj, 2022; Liu et al., 2022a; Ye et al., 2021). However, flares in their dataset are unrealistic.
@@
There has been active research on UDC image restoration. DISCNet (Feng et al., 2021) incorporates the domain knowledge of the image formation model. UDC-UNet (Liu et al., 2022b), a second performer of MIPI challenge (Feng et al., 2022), introduces kernel branches to incorporate prior knowledge and condition branches for spatially variant manipulation.
@@
Many studies have focused on video restoration models for general tasks, such as denoising (Tassano et al., 2020), deblurring (Wang et al., 2019; Zhong et al., 2020), and super-resolution (Wang et al., 2019). Unlike image restoration, which only focuses on a spatial dimension, video restoration leverages temporal information. FastDVDNet (Tassano et al., 2020) uses a two-step denoising process in a multi-scale architecture to leverage temporal information without explicit motion estimation. EDVR (Wang et al., 2019) aligns features using deformable convolutions (Dai et al., 2017) and applies both temporal and spatial attention to highlight essential features. ESTRNN (Zhong et al., 2020) integrates residual dense blocks into RNN cells for spatial feature extraction and employs a spatiotemporal attention module for feature fusion. However, studies on UDC video restoration are still rare. DDRNet (Liu et al., 2024), the pioneering work to address UDC video degradation, adopts a recurrent architecture that merges multi-scale feature learning and bi-directional propagation.
@@ 3 Dataset Acquisition
Since obtaining well-synchronized and precisely aligned paired videos for the same scene is challenging, we carefully design both hardware and software for capturing videos.
@@
3.1 The Video Capturing System
As shown in Figure 2, we present a UDC video capturing system consisting of two camera modules, a display panel for the UDC area, a beam splitter, two 6-axis stages, and a single-board computer. In this setup, one of the two camera modules is under low light conditions caused by the display panel, making synchronization between paired frames more challenging than in previous beam splitter setups (Hwang et al., 2015; Joze et al., 2020; Li et al., 2023; Rim et al., 2020). To capture synchronized videos for the same scene, we propose a UDC video-capturing system that ensures precise camera synchronization and accurate frame alignment.
@@ 3.1 The Video Capturing System
We use the Hawk-Eye (IMX686) (Arducam, 2022) to ensure that UDC-VIT exhibits a similar UDC degradation as Samsung Galaxy Z-Fold 5’s UDC. Both devices use Quad Bayer Coding (QBC), a technique designed to mitigate the reduced sensor luminance sensitivity often associated with higher camera resolutions or smaller sensor pixel sizes (Sony, 2014; Wikipedia, 2014). As shown in Figure 3, four adjacent pixels share the same color filter in the quad Bayer structure. These pixels are grouped to increase sensitivity and reduce noise in low-light conditions (e.g., in the UDC setting). Conversely, in bright conditions, the sensor reverts the pixels to the Bayer structure through the remosaicing process, maintaining the Bayer sensor’s high resolution.
@@
We use a non-polarizing cube-shaped beam splitter (e.g., Thorlabs CCM1-BS013 (Thorlabs, 2015)) to enable the two camera modules to capture the same scene. The beam splitter divides the incident light into two directions with a beam deviation of $0\pm 5$ arcminutes at a 50:50 ratio. The unused optical path of the combiner is black-coated to minimize image contrast loss caused by scattering. They can capture the same scene by aligning the two cameras to the beam splitter’s split fields of view (FOV). Figure 2(a) illustrates the optomechanical layout of the dual camera and beam splitter.
@@
Despite many studies using beam splitters for paired image dataset collection (Hwang et al., 2015; Joze et al., 2020; Li et al., 2023; Rim et al., 2020), this paper is the first to apply them in UDC research to the best of our knowledge. It presents challenges to align the optical paths of the display panel’s UDC area, a beam splitter, and two camera modules. To ensure alignment between the cameras’ optical axes and the beam splitter, we employ Thorlabs K6XS 6-axis kinematic optical mounts (Thorlabs, 2013). Each camera module is mounted on a K6XS mount, allowing for shifts, rotations, and tilts across the six axes to align their FOV.
@@
We use a Raspberry Pi 5 (Arducam, 2023) that has two four-lane MIPI interface connections for high bandwidth to synchronize the two high-resolution cameras. It ensures stable high-resolution video recording. To synchronize the cameras, we use independent streamers managed by MPI barriers (Message Passing Interface Forum, 2023), achieving synchronization with an accuracy margin of up to 8 msec. Consistent frame rates for both cameras are ensured using the uncompressed binary dump method (e.g., YUV420 format). Despite these settings resulting in less than a 0.5 fps (8 msec) difference between paired frames, rapid movements may still cause the cameras to capture different scenes. Thus, videos capturing fast-moving objects (e.g., speeding cars) are excluded from the dataset.
@@ 3.2 Obtaining Aligned Video Pairs
This section illustrates how we align the optical axis and FOV, the criteria for determining FOV alignment, and the test cases. We use a real-time viewing system for the two cameras. We roughly align the view, fine-tune it using the K6XS and DFT alignment, evaluate the accuracy evaluation (e.g., PCK), record the video, and make the final selection manually. Please see Section 4 for detailed information on the PCK.
Algorithm 1 Aligned video capturing process for UDC-VIT.
The alignment accuracy of the paired frames is greater than 90%.
while The average PCK ($\alpha=0.005$) $<$ 90% do
     Initial setup. Adjust the camera positions and the beam splitter, ensuring that the views of the two cameras are roughly similar.
     Fine-tuning. The rotation, tilt, and horizontal/vertical positions of the K6XS are finely adjusted by observing a $12\times 9$ checkerboard and everyday scenes in the live view system.
     DFT alignment and PCK evaluation. Align paired frames using DFT and calculate the average PCK.
@@
Final selection. Only the videos all authors assessed aligned and synchronized are retained.
@@
Despite the careful design of our video-capturing system, unavoidable misalignments, such as shifts, rotations, and tilts, still occur between paired frames. Previous methods, such as SIFT (Lowe, 2004), RANSAC (Fischler & Bolles, 1981), and deep learning approaches (Feng et al., 2023), struggle to perform well in the existence of severe degradation introduced by the UDC. Thus, we use DFT to achieve degradation-resilient alignment (Ahn et al., 2024).
The alignment process is summarized as shift, rotate, and crop paired frames using DFT. Captured videos have an original frame size of $(1920,1080,3)$. The ground-truth frame is center-cropped to $(1900,1060,3)$, and the degraded frame undergoes a cropping around the center. To align the cropped degraded frame $\mathcal{D}$ with the cropped ground-truth frame $\mathcal{G}$, we iteratively shift the $(x,y)$ coordinates and rotate the frames to find the point of minimum loss. Our focus is on addressing shifts and rotations while excluding tilts. Handling tilts is challenging because of the need for perspective transforms optimized for objects in the same plane within a single image. Despite not considering tilts, our video-capturing system minimizes all shifts, rotations, and tilts so that they do not significantly affect the alignment, as confirmed by our experiment (the PCK values in Table 2). The loss function $\mathcal{L}$ for the alignment between $\mathcal{D}$ and $\mathcal{G}$ is defined as below:
$\mathcal{L}=\lambda_{1}\sum\limits_{x=0}^{M-1}\sum\limits_{y=0}^{N-1}(\mathcal{D}(x,y)-\mathcal{G}(x,y))^{2}+\lambda_{2}\sum\limits_{u=0}^{M-1}\sum\limits_{v=0}^{N-1}\Delta\mathcal{F}_{amp}(u,v)+\lambda_{3}\sum\limits_{u=0}^{M-1}\sum\limits_{v=0}^{N-1}\Delta\phi(u,v),$ (1)
where the first term is the mean squared error, $\Delta\mathcal{F}_{amp}(u,v)$ and $\Delta\phi(u,v)$ represent the L1 distance for the amplitude and phase, respectively. They are defined as $\Delta\mathcal{F}_{amp}(u,v)=|\mathcal{F}_{\mathcal{D}}(u,v)-\mathcal{F}_{\mathcal{G}}(u,v)|$ and $\Delta\phi(u,v)=|\phi_{\mathcal{D}}(u,v)-\phi_{\mathcal{G}}(u,v)|$. Note that $\mathcal{F}(u,v)$ is the frequency value at the point $(u,v)$ in the frequency domain. Following Ahn et al.’s setting, we use $\lambda_{1}=\lambda_{3}=1,\lambda_{2}=0$. The detailed alignment algorithm is described in Appendix A.

### v2

@@ Abstract
Even though an Under-Display Camera (UDC) is an advanced imaging system, the display panel significantly degrades captured images or videos, introducing low transmittance, blur, noise, and flare issues. Tackling such issues is challenging because of the complex degradation of UDCs, including diverse flare patterns. However, no dataset contains videos of real-world UDC degradation. In this paper, we propose a real-world UDC video dataset called UDC-VIT. Unlike existing datasets, UDC-VIT exclusively includes human motions for facial recognition. We propose a video-capturing system to acquire clean and UDC-degraded videos of the same scene simultaneously. Then, we align a pair of captured videos frame by frame, using discrete Fourier transform (DFT). We compare UDC-VIT with six representative UDC still image datasets and two existing UDC video datasets. Using six deep-learning models, we compare UDC-VIT and an existing synthetic UDC video dataset. The results indicate the ineffectiveness of models trained on earlier synthetic UDC video datasets, as they do not reflect the actual characteristics of UDC-degraded videos. We also demonstrate the importance of effective UDC restoration by evaluating face recognition accuracy concerning PSNR, SSIM, and LPIPS scores. UDC-VIT is available at our official GitHub repository.
@@ 1 Introduction
Under-Display Camera (UDC) is an imaging system where the camera is positioned beneath the display. Modern smartphones, including the Samsung Galaxy Z-Fold series [43, 44, 45, 46] and the ZTE Axon series [64, 65, 66] have adopted UDCs. The UDC area, depicted in Fig. 1, serves as display space under normal circumstances and acts as the light’s passage to the camera when capturing pictures or videos. This design allows for a larger screen-to-body ratio, meeting the common consumer demand for a full-screen display without a camera hole or notch. However, UDC introduces severe and complex degradations such as reduced transmittance, noise, blur, and flare in a single image or video frame. Moreover, motion is also involved in UDC videos.
@@
UDC degradation results from micrometer-scale diffraction of incoming light by display pixels [39]. To mitigate this, modern UDC smartphones adopt lower pixel density in the UDC area, as shown in Fig. 1(c). However, this reduces the local display resolution and disrupts the natural video viewing experience, making high-quality video restoration essential.
@@
Many studies have explored UDC image datasets, including synthetic datasets (e.g., T-OLED/P-OLED [63] and SYNTH [18]) and a pseudo-real dataset [20]. Additionally, a real-world dataset called UDC-SIT [1] exists.
@@
Ahn et al[1] highlight the importance of training DNN models on real-world UDC datasets as synthetic datasets do not reflect the actual characteristics of UDC-degraded images. However, a real-world UDC video dataset and restoration model have yet to be introduced. Although several studies address the synthetic UDC video datasets [12, 31], they still fall short of accurately representing the actual properties of UDC videos. There are two main challenges in constructing a real-world UDC video dataset. One is to find a matching pair of UDC-distorted and ground-truth videos with high alignment accuracy. The other is to synchronize the time for all frames when capturing videos.
@@
This paper proposes a new UDC video dataset called UDC-VIT (UDC’s VIdeo by Thunder Research Group). As far as we know, it is the first real-world UDC video dataset to address limitations of the existing UDC video datasets.
@@
Using a non-polarizing cube beam splitter [56], we create a video-capturing system to minimize discrepancies between paired frames. We cut the UDC area of a smartphone display (e.g., Samsung Galaxy Z-Fold 5 [45]) and attach it to the beam splitter. Two Arducam Hawk-Eye (IMX686) camera modules [7] are placed on both sides of the beam splitter. These modules, operated by a Raspberry Pi 5 [8], capture synchronized video frame pairs using the Message Passing Interface (MPI) barrier.
Fig. 2 shows our UDC video capturing system. Despite the meticulous design, inevitable pixel-position difference occurs. We correct this difference between the two matched frames for the same scene by using the DFT [11] following the previous work by Ahn et al[1].
The contributions of this paper are as follows:
  • We address the limitations of existing datasets, including unrealistic degradations, flares, and artifacts, emphasizing the need for a real-world dataset (Sec. 4 and Appendix A).
  • We provide UDC-VIT, a real-world UDC video dataset that accurately reflects actual UDC degradations, ensuring precise spatial and temporal alignment through our meticulously designed video-capturing system (Sec. 3 and Appendix D).
  • We describe UDC-VIT’s effectiveness through extensive experiments, comparing it with an existing synthetic dataset using six deep-learning models (Sec. 5, Appendix B, and Appendix C).
  • We highlight the importance of restoring UDC degradation for applications like Face ID by evaluating face recognition accuracy at different restoration levels. Our dataset includes real-world face videos, enhancing its relevance for practical use (Sec. 5).
@@ 2 Related Work
There has been extensive research on UDC still image datasets. Zhou et al[63] propose the T-OLED/P-OLED datasets. Images are displayed on a monitor, and paired images are captured with and without a T-OLED/P-OLED display in front of the camera. However, due to the limited dynamic range of the monitor, flares are almost absent in their datasets. Feng et al[18] propose the SYNTH dataset. They convolve the measured point spread function (PSF) of ZTE Axon 20 [64] with clean images [24], exhibiting flares. However, it has limitations such as the absence of noise and spatially variant flares. Notably, UDC distortion gradually increases from the center of the camera lens to outwards, leading to spatially distorted flares [61]. Feng et al[20] propose a pseudo-real dataset by capturing paired images of similar scenes using two cameras (e.g., ZTE Axon 20 UDC [64] and iPhone 13 Pro camera [6]). However, they use two cameras, leading to geometric misalignment. They improve the geometric misalignment using AlignFormer [20]. Nonetheless, they encounter challenges with alignment accuracy. Ahn et al. propose a real-world dataset called UDC-SIT and an image-capturing system [1, 4]. They attach Samsung Galaxy Z-Fold 3 [43]’s UDC area to a lid. Paired images are acquired by opening and closing the lid onto the Samsung Galaxy Note 10’s standard camera [42]. They use DFT to align the misalignment between the paired images that occurs during the opening and closing of the lid [5]. The images in the UDC-SIT dataset contain the actual UDC degradation (e.g., spatially variant flares). Finally, Wang et al[58] and Tan et al[53] propose face datasets for facial expression recognition or restoration. They are synthesized using a GAN-based model trained on the T-OLED/P-OLED datasets [63], which lacks realistic UDC degradation, particularly flares. Moreover, the datasets are not publicly available.
@@
Research has been conducted on synthetic UDC video datasets. Chen et al[12] propose the PexelsUDC-T/P datasets. They train a GAN-based UDC video generation model using T-OLED/P-OLED datasets [63], which do not show flares. They generate UDC-degraded videos using clean videos [38]. The datasets are not publicly available. Liu et al[31] propose the VidUDC33K dataset. They convolve the measured PSF on the clean video frames [24] to show flares. They simulate the dynamic change of the PSF [29] between consecutive frames following the previous work [9, 32, 60]. However, flares in their dataset are unrealistic.
@@
There has been active research on UDC image restoration. DISCNet [18] incorporates the domain knowledge of the image formation model. UDC-UNet [33], a second performer of MIPI challenge [19], introduces kernel branches to incorporate prior knowledge and condition branches for spatially variant manipulation. SFIM [2] integrates spatial and frequency information to effectively address flare artifacts in real UDC images.
@@
Many studies have focused on video restoration models for general tasks, such as denoising [54], deblurring [57, 62], and super-resolution [57]. Unlike image restoration, which only focuses on a spatial dimension, video restoration leverages temporal information. FastDVDNet [54] uses a two-step denoising process in a multi-scale architecture to leverage temporal information without explicit motion estimation. EDVR [57] aligns features using deformable convolutions [13] and applies both temporal and spatial attention to highlight essential features. ESTRNN [62] integrates residual dense blocks into RNN cells for spatial feature extraction and employs a spatiotemporal attention module for feature fusion. However, studies on UDC video restoration are still rare. DDRNet [31], the pioneering work to address UDC video degradation, adopts a recurrent architecture that merges multi-scale feature learning and bi-directional propagation.
@@ 3 Dataset Acquisition
Since obtaining well-synchronized and precisely aligned paired videos for the same scene is challenging, we carefully design both hardware and software for video capture.
3.1 The Video Capturing System
As shown in Fig. 2, we present a UDC video capturing system with two camera modules, a display panel for the UDC area, a beam splitter, two 6-axis stages, and a single-board computer. In this setup, one of the two camera modules is in low light conditions caused by the display panel, making synchronization between paired frames more challenging than in previous beam splitter setups [25, 26, 30, 41]. To capture synchronized videos for the same scene, we propose a UDC video-capturing system that ensures precise camera synchronization and frame alignment.
@@ 3.1 The Video Capturing System
We use the Hawk-Eye (IMX686) [7] to ensure that UDC-VIT exhibits a similar UDC degradation as Samsung Galaxy Z-Fold 5’s UDC. Both devices use Quad Bayer Coding (QBC), a technique designed to mitigate the reduced sensor luminance sensitivity often associated with higher camera resolutions or smaller sensor pixel sizes [49, 59]. As shown in Fig. 3, four adjacent pixels share the same color filter in the quad Bayer structure. These pixels are grouped to increase sensitivity and reduce noise in low-light conditions (e.g., in the UDC setting). Conversely, in bright conditions, the sensor reverts the pixels to the Bayer structure through the remosaicing process, maintaining the Bayer sensor’s high resolution. We apply pixel binning and ISP to both GT and degraded captures to ensure resolution consistency and synchronization via stable auto-exposure in well-lit conditions. Please see Appendix D for details.
@@
We use a non-polarizing cube-shaped beam splitter (e.g., Thorlabs CCM1-BS013 [56]) to enable the two cameras to capture the same scene. The beam splitter divides the incident light into two directions with a beam deviation of 0±50$\pm$ 50 ± 5 arcminutes at a 50:50 ratio. The unused optical path is black-coated to minimize image contrast loss due to scattering. They can capture the same scene by aligning the two cameras to the beam splitter’s split fields of view (FOV). Fig. 2(a) illustrates the optomechanical configuration of our beam splitter–based dual-camera setup.
@@
Despite many studies using beam splitters for paired image dataset collection [25, 26, 30, 41], this paper is the first to apply them in UDC research. It presents challenges to align the optical paths of the display panel’s UDC area, a beam splitter, and two camera modules. To ensure alignment between the cameras’ optical axes and the beam splitter, we employ Thorlabs K6XS 6-axis kinematic optical mounts [55]. Each camera module is mounted on a K6XS mount, allowing for shifts, rotations, and tilts across the six axes to align their FOV.
@@
We use a Raspberry Pi 5 [8] that has two four-lane MIPI interface connections for high bandwidth to synchronize the two high-resolution cameras. It ensures stable high-resolution video recording. To synchronize the cameras, we use independent streamers managed by MPI barriers [35], achieving synchronization with an accuracy margin of up to 8 msec. Consistent frame rates for both cameras are ensured using the uncompressed binary dumps (e.g., YUV420 format). Despite these settings resulting in less than a 0.5 fps (8 msec) difference between paired frames, rapid movements may still cause the cameras to capture different scenes. Thus, videos capturing such objects (e.g., speeding cars) are excluded from the dataset.
@@ 3.2 Obtaining Aligned Video Pairs
This section explains optical axis and FOV alignment, criteria for FOV alignment, and test cases. As illustrated in Algorithm 1, using a real-time viewing system, we roughly align the cameras, fine-tune them with K6XS and DFT, evaluate accuracy (e.g., PCK), record the video, and manually select the final footage. For PCK details, see Sec. 4.
Algorithm 1 Aligned video capturing process.
Paired frame alignment accuracy exceeds 90%.
while Average PCK ($\alpha=0.005$) $<$ 90% do
  Initial setup. Adjust camera positions and beam splitter for roughly similar views.
  Fine-tuning. Precisely adjust rotation/tilt angles and horizon/vertical position of K6XS by observing a 12×912$\times$ 912 × 9 checkerboard and everyday scenes in the live view system.
  Alignment and evaluation. Align paired frames using DFT and compute the average PCK.
@@
Final selection. Retain only aligned and synchronized videos approved by all authors.
@@
Despite the careful design of our video-capturing system, misalignments, such as shifts, rotations, and tilts, still occur between paired frames. Previous alignment methods, such as SIFT [34], RANSAC [21], and deep learning approaches [20], struggle to perform well in the existence of severe degradation introduced by the UDC. Thus, we use DFT to achieve degradation-resilient alignment [1]. Please see Appendix D for details.
[…]
$\begin{split}\mathcal{L}=&\lambda_{1}\sum\limits_{x=0}^{M-1}\sum\limits_{y=0}^{N-1}(\mathcal{D}(x,y)-\mathcal{G}(x,y))^{2}\\ &+\lambda_{2}\sum\limits_{u=0}^{M-1}\sum\limits_{v=0}^{N-1}\Delta\mathcal{F}_{amp}(u,v)+\lambda_{3}\sum\limits_{u=0}^{M-1}\sum\limits_{v=0}^{N-1}\Delta\phi(u,v),\end{split}$ (1)
[…]